Solve for what is being asked.



What is the molar mass of the solute in a solution made by dissolving 34.41



g of an unknown nonelectrolyte compound in 280.0 g benzene. The boiling point of



the solution is 93.2 °C. Benzene has a boiling point of 80.1 °C and molal boiling-



point elevation constant of 2.53 °C/m.

1
Expert's answer
2022-04-18T17:03:09-0400

m(solvent, benzene) = 280.0 g = 0.2800 kg;

m(unknown nonelectrolyte) = 34.41 g;

Kb = 2.53 C/m;

T = 93.2 C

T0 = 80.1 C;

"\Delta" T = T - T0;

"\Delta" T = Kb * m = Kb * n/m(solvent) = Kb * m(unknown nonelectrolyte)/(M(unknown nonelectrolyte) * m(solvent);

M(unknown nonelectrolyte) = Kb * m(unknown nonelectrolyte)/("\Delta" T * m(solvent)) = Kb * m(unknown nonelectrolyte)/((T - T0) * m(solvent)) = 2.53 * 34.41/((93.2-80.1) * 0.2800) = 2.53 * 34.41/(13.1 * 0.2800) = 23.7 g/mol.

Answer: 23.7 g/mol.